# Convergence of the generalization error for Deep Gradient Flow Methods for PDEs

### v1

@@ 1. Introduction
Deep learning methods for the solution of high-dimensional partial differential equations (PDEs) have gained tremendous popularity in the last few years, since they can tackle equations in dimensions that were not attainable by classical methods, such as finite difference and finite element schemes. This ability allows the modeling of more realistic phenomena across various fields of science and technology, including engineering, biology, economics, and finance. The seminal articles of Sirignano and Spiliopoulos [29] on the Deep Galerkin Method (DGM) and of Raissi and Karniadakis [26] and Raissi et al. [27] on physics-informed neural networks (PINNs), building on the earlier work of Lagaris et al. [18] and Lagaris et al. [19], incorporate the PDE residual and the initial and boundary conditions into the loss function of a neural network, which is then minimized by stochastic gradient descent (SGD). These methods have laid the foundations for a variety of extensions and applications, including among many others, fractional differential equations (Pang et al. [23]), variational PINNs (Kharazmi et al. [17]), Bayesian variants (Yang et al. [31]) and mean-field games (Carmona and Zeng [4]).
On the other hand, deep gradient flow methods (DGFMs), also known as deep Ritz methods, formulate the PDE as an energy minimization problem, where the energy is derived from the differential operator, which typically leads to a loss function that is easier to compute. Moreover, they usually discretize the equation in time and train one network for each time step, instead of using a monolithic space-time discretization; see e.g., E and Yu [6], Liao and Ming [20], Georgoulis et al. [8], Park et al. [25], Papapantoleon and Rou [24], Bruna et al. [3] for differential operators, and Georgoulis et al. [9] for an integro-differential operator. A comprehensive review of the available methods appears in the forthcoming book of Jentzen et al. [14].
@@
where $\mathcal{A}$ is a differential operator, $\Phi$ determines the initial condition, $T$ is a (finite) time horizon, and $D\subseteq\mathbb{R}^{d}$ is the domain of the PDE. The DGFMs translate the PDE into an energy minimization problem, which is then computed using stochastic gradient descent or one of its variants (e.g. ADAM), and can be described in the following manner:
@@
where $\mathcal{C}^{n}_{\theta}$ denotes the space of neural networks with $n$ neurons where $\theta$ is the set of trainable parameters, while $\ell$ denotes the energy functional associated to the operator $\mathcal{A}$.
@@
The generalization error $\mathcal{E}_{\text{gen}}$ can be decomposed in three separate components:
@@
  • the training error $\mathcal{E}_{\text{train}}$, which refers to how well GD or SGD approximate the true solution of the minimization problem (1.2).
@@
$\mathcal{E}_{\text{gen}}=\mathcal{E}_{\text{train}}+\mathcal{E}_{\text{quad}}+\mathcal{E}_{\text{approx}},$ (1.3)
@@
and the aim of the present paper is to study these errors and show that, as the numbers of neurons tends to infinity and the training time also tends to infinity, then the generalization error tends to zero, and the outcome of the deep gradient flow method indeed approximates the solution of the PDE.
@@
There are several articles available that study the generalization error of deep learning methods for PDEs, typically focusing on the popular DGM and PINN methods. These methods rely on approximability properties of neural networks and properties of quadrature methods in order to control the generalization error, while they typically consider only a posteriori estimates for the training error. We refer the interested reader to Mishra and Molinaro [22] and Gazoulis et al. [7] for results on PINNs, and the related while more general article of Loulakis and Makridakis [21]. Moreover, several articles consider the approximation error of DMG and PINNs; see, for example, Sirignano and Spiliopoulos [29] and Shin et al. [28]. The recent article of Jiang et al. [15] considers the “global” convergence of DGM and PINNs, which amounts to the convergence of the training error in our notation. Combined with other available results, this article allows to deduce the convergence of the generalization error of these methods. Compared to the extended literature on DGM and PINNs, there are significantly fewer papers on DGFMs; let us mention here the articles of Dondl et al. [5] which focuses on the approximation error, and Jiao et al. [16] which provides a convergence rate using the Rademacher complexities.
@@
The aim of the present article is to provide convergence results on the generalization error of DGFMs under reasonable and verifiable hypothesis on the underlying PDEs. The first part of this work focuses on the analysis of the approximation error, i.e. we show that there exists a neural network that approximates the solution of the PDE. This result uses ideas from PDE theory, optimization and the calculus of variations, and is inspired by the seminal paper of Sirignano and Spiliopoulos [29]. The second part of this work focuses on the analysis of the training error, and we show that as the number of neurons tend to infinity and the training time also tends to infinity, then the outcome of the deep gradient flow method tends to the true solution of the PDE. This result is inspired by the work of Jiang et al. [15]. The quadrature error is the most well-understood error of the three, thus this work focuses on the other two errors. Moreover, let us mention for the sake of completeness, that our method also induces a discretization error, from the time-stepping scheme. However, as this error is also well-studied and understood, we have chosen to omit it from the discussion here. The combination of these results, yields that the generalization error also tends to zero.
@@
This article is organized as follows: 2 provides an overview of deep gradient flow methods for the solution of PDEs. 3 studies the approximation error of DGFMs, using the variational formulation of PDEs and a tailored version of the universal approximation theorem. 4 studies the training error of DGFMs; we first derive a gradient flow that the training process satisfies in the “wide network limit” and then analyze the behavior of this flow as the training time tends to infinity. Finally, the appendices contain auxilliary estimates and examples.
@@ 2. Deep gradient flow methods for PDEs
Let us start by providing an overview of deep gradient flow methods (DGFMs) for the solution of PDEs. These methods have gained increased popularity in the literature because they can efficiently handle high-dimensional PDEs stemming from physics, engineering, and finance; see e.g. E and Yu [6], Liao and Ming [20], Georgoulis et al. [8], Park et al. [25] and Papapantoleon and Rou [24] for differential operators, and Georgoulis et al. [9] for an integro-differential operator. Deep gradient flow methods reformulate the PDE as an energy minimization problem, which is then approximated in a time-stepping fashion by deep artificial neural networks. This method results in a loss function that is tailor-made to the PDE at hand, avoids the use of a second derivative, which is computationally costly, and reduces the training time compared to, for instance, the DGM of Sirignano and Spiliopoulos [29]; see e.g. Georgoulis et al. [8, Sec. 5].
@@
where $\mathcal{L}$ is a self-adjoint, linear operator and $F$ is a (possibly non-linear) operator from $\mathcal{V}$ to $\mathcal{V}^{*}$. This PDE is then discretized using, for example, the backward Euler differentiation scheme, which yields
@@
Next, we present examples of PDEs that have been treated by DGFMs, and their applications.
@@ 3.1. Well-posedness
According to Hilber et al. [12, Theorem 3.2.2], we only need to verify that the bilinear form $\left\langle\mathcal{A}u,v\right\rangle_{\mathcal{H}^{-1},\mathcal{H}_{0}^{1}}$ is continuous and satisfies the “Gårding inequality”, where
@@ 3.2. Time stepping
The proof follows directly from Theorem 2.1 in Akrivis et al. [1]. Indeed, using that $U^{0}=u\left(0\right)$, we can show by direct, but tedious, calculations that the assumptions of [1, p. 523] are satisfied for $\lambda<1$ and $q=1$. ∎
@@ 3.4. Neural network approximation and a version of the Universal Approximation Theorem
Observe that $\mathcal{C}\left(\psi\right)\subset\mathcal{H}_{0}^{1}(\mathbb{R}^{d})$, since $\psi\in C^{\infty}_{c}(\mathbb{R}^{d})$. Assume that $\mathcal{C}\left(\psi\right)$ is not dense in $\mathcal{H}_{0}^{1}(\mathbb{R}^{d})$ then, as a corollary of the Hahn–Banach extension theorem, see e.g. van Neerven [30, Corollary 4.12], there exists a non-zero continuous linear functional $G$ on $\mathcal{H}_{0}^{1}(\mathbb{R}^{d})$ such that for any $\zeta\in\mathcal{C}\left(\psi\right)$,
@@
Since $g^{\varepsilon}_{1}-\nabla\cdot\left(g^{\varepsilon}_{2}\right)$ is continuous, see e.g. van Neerven [30, Proposition 11.1], by Lemma 3.12, we conclude that $g^{\varepsilon}_{1}-\nabla\cdot\left(g^{\varepsilon}_{2}\right)=0$. Then, for any $f\in C^{\infty}_{c}(\mathbb{R}^{d})$,
@@ 3.5. Convergence of the minimizer
Next, we prove that $I^{k}\left(w_{m}\right)\to I^{k}\left(w_{*}\right)$ implies that $w_{m}\to w_{*}$ in $\mathcal{H}_{0}^{1}$. Let us first notice that $w_{m}\rightharpoonup w_{*}$. Otherwise, there exists a subsequence $(w_{m_{i}})$, an ${\varepsilon}>0$ and a nonzero functional $f$ such that $\left|f\left[w_{m_{i}}\right]-f\left[w_{*}\right]\right|\geq{\varepsilon}$. Since $(w_{m_{i}})$ is bounded in $\mathcal{H}_{0}^{1}$ (otherwise $I^{k}\left(w_{m_{i}}\right)$ is unbounded), it is pre-weakly compact (which means it has a weakly convergent subsequence, see e.g. van Neerven [30, Corollary 4.56]). Let us denote one of its weak limits by $w_{**}$. Using Lemma 3.5, $I^{k}\left(w_{*}\right)=\lim_{i\to\infty}I^{k}\left(w_{m_{i}}\right)\geq I^{k}\left(w_{**}\right)$. This inequality implies $w_{*}=w_{**}$ by the uniqueness of the minimizer. Hence $w_{m}\rightharpoonup w_{*}$, a contradiction with $\left|f\left[w_{m_{i}}\right]-f\left[w_{*}\right]\right|\geq{\varepsilon}$. Therefore, $w_{m}\rightharpoonup w_{*}$.
@@ 4.1. Convergence of the trained neural network
for some $r_{n}$ increasing with $n$. We restrict the domain of the parameters $(\beta^{i},\alpha^{i},c^{i})$ to $r_{n},r_{n}]$ which converges to $\mathbb{R}$ as $n\to\infty$, and for $\alpha^{i}$ we also need to subtract the ball $\left(-\frac{1}{r_{n}},\frac{1}{r_{n}}\right)$. Gradient clipping is in accordance with deep learning literature, see, for example, Zhang et al. [32] and Goodfellow et al. [10, Ch. 10 and 11].
@@ 4.2. Long time behavior of the gradient flow
Using van Neerven [30, Proposition 14.13], every trace class operator is compact and positive definite. Therefore, we can do a spectral decomposition for the operator $\widetilde{\mathcal{T}}$. There exists an orthogonal basis $\{\tilde{\mathrm{e}}_{i}\}^{\infty}_{i=1}$, such that

### v2

@@ 1. Introduction
Deep learning methods for the solution of high-dimensional partial differential equations (PDEs) have gained tremendous popularity in the last few years, since they can tackle equations in dimensions that were not attainable by classical methods, such as finite difference and finite element schemes. This ability allows the modeling of more realistic phenomena across various fields of science and technology, including engineering, biology, economics, and finance. The seminal articles of Sirignano and Spiliopoulos [37] on the Deep Galerkin Method (DGM) and of Raissi and Karniadakis [32] and Raissi et al. [33] on physics-informed neural networks (PINNs), building on the earlier work of Lagaris et al. [23] and Lagaris et al. [24], incorporate the PDE residual and the initial and boundary conditions into the loss function of a neural network. This loss function is then minimized by stochastic gradient descent (SGD) or one of its variants, e.g. AdaGrad, Adam or RMSProp; see, for example Latz [25], Jin et al. [20, 21] and Sirignano and Spiliopoulos [36, 38] for more details on SGD. These methods have laid the foundations for a variety of extensions and applications, including among many others, fractional differential equations (Pang et al. [29]), variational PINNs (Kharazmi et al. [22]), Bayesian variants (Yang et al. [40], Ding et al. [6]) and mean-field games (Carmona and Zeng [5]); see also Grossmann et al. [13] and Sikora et al. [35] for a comparison of PINNs with finite element methods.
On the other hand, deep gradient flow methods (DGFMs), also known as deep Ritz methods, formulate the PDE as an energy minimization problem, where the (Dirichlet) energy is derived from the differential operator, which typically leads to a loss function that is easier to compute. Moreover, they usually discretize the equation in time and train one network for each time step, instead of using a monolithic space-time discretization; see e.g., E and Yu [8], Liao and Ming [26], Georgoulis et al. [10], Park et al. [31], Papapantoleon and Rou [30], Bruna et al. [4] for differential operators, and Georgoulis et al. [11] for an integro-differential operator. A comprehensive review of deep learning methods for the solution of PDEs (and related BSDEs) appears in the forthcoming book of Jentzen et al. [17].
@@
where $\mathcal{A}$ is a differential operator, $\Phi$ determines the initial condition, $T$ is a (finite) time horizon, and $D\subseteq\mathbb{R}^{d}$ is the domain of the PDE. DGFMs translate the PDE into an energy minimization problem, which is then computed using stochastic gradient descent or one of its variants. The solution computed by DGFMs can be described in the following manner:
@@
where $\mathcal{C}^{n}_{\theta}$ denotes the space of neural networks with $n$ neurons where $\theta$ is the set of trainable parameters, while $\ell$ denotes the loss functional associated to the Dirichlet energy of the operator $\mathcal{A}$.
@@
The generalization error $\mathcal{E}_{\text{gen}}$ can be decomposed in four separate components:
@@
  • the time-stepping error $\mathcal{E}_{\text{step}}$, which refers to how well the time-stepping scheme is approximating the true PDE;
@@
  • the training error $\mathcal{E}_{\text{train}}$, which refers to how well SGD approximates the true solution of the minimization problem (1.2).
@@
$\mathcal{E}_{\text{gen}}=\mathcal{E}_{\text{quad}}+\mathcal{E}_{\text{step}}+\mathcal{E}_{\text{approx}}+\mathcal{E}_{\text{train}},$ (1.3)
@@
and the aim of the present paper is to study these errors and show that, as the number of quadrature points, time-steps, neurons and training time tends to infinity, then the generalization error tends to zero, and the outcome of the deep gradient flow method indeed approximates the solution of the PDE.
@@
There are several articles available that study the generalization error of deep learning methods for PDEs, typically focusing on the popular DGM and PINN methods. These methods rely on approximability properties of neural networks and properties of quadrature methods in order to control the generalization error, while they typically consider only a posteriori estimates for the training error. We refer the interested reader to Mishra and Molinaro [28] and Gazoulis et al. [9] for results on PINNs, and the related while more general article of Loulakis and Makridakis [27]. Moreover, several articles consider the approximation error of DMG and PINNs; see, for example, Sirignano and Spiliopoulos [37], Shin et al. [34] and Abdo et al. [1]. The recent article of Jiang et al. [18] considers the “global” convergence of DGM and PINNs, which amounts to the convergence of the training error in our notation. Combined with other available results, this article allows to deduce the convergence of the generalization error of the DGM and PINN methods. Compared to the extensive literature on DGM and PINNs, there are significantly fewer papers on the convergence analysis of DGFMs; let us mention here the articles of Dondl et al. [7] which focuses on the approximation error, and Jiao et al. [19] which provides a convergence rate using the Rademacher complexities.
@@
The aim of the present article is to provide convergence results on the generalization error of DGFMs under reasonable and verifiable hypotheses on the underlying PDEs. The first part of this work focuses on the analysis of the approximation error, i.e. we show that there exists a neural network that approximates the solution of the PDE. This result uses ideas from PDE theory, optimization and the calculus of variations, and a suitable variant of the universal approximation theorem for neural networks. This part is inspired by the seminal paper of Sirignano and Spiliopoulos [37], although the methods are rather different. The second part of this work focuses on the analysis of the training error, and we show that as the number of neurons tend to infinity and the training time also tends to infinity, then the outcome of the deep gradient flow method tends to the true solution of the PDE. This result is based on the careful analysis of the gradient flow associated with the loss function of DGFMs. This part follows the work of Jiang et al. [18], utilizing the structure of DGFMs. The quadrature error and the discretization error of the time-stepping scheme are the most well-understood errors in the error decomposition (1.3), thus this work focuses on the other two errors. The combination of these results, yields that the generalization error of DGFMs, under reasonable and verifiable assumptions, tends to zero.
@@
This article is organized as follows: Section 2 provides an overview of deep gradient flow methods for the solution of PDEs. Section 3 studies the approximation error of DGFMs, using the variational formulation of PDEs and a tailored version of the universal approximation theorem. Section 4 studies the training error of DGFMs; we first derive a gradient flow that the training process satisfies in the “wide network limit” and then analyze the behavior of this flow as the training time tends to infinity. Finally, the appendices contain auxiliary estimates and examples.
@@ 2. Deep gradient flow methods for PDEs
Let us start by providing an overview of deep gradient flow methods (DGFMs) for the solution of PDEs. These methods have gained increased popularity in the literature because they can efficiently handle high-dimensional PDEs stemming from physics, engineering, and finance; see e.g. E and Yu [8], Liao and Ming [26], Georgoulis et al. [10], Park et al. [31] and Papapantoleon and Rou [30] for differential operators, and Georgoulis et al. [11] for an integro-differential operator. Deep gradient flow methods reformulate the PDE as an energy minimization problem, which is then approximated in a time-stepping fashion by deep artificial neural networks. This method results in a loss function that is tailor-made to the PDE at hand, avoids the use of a second derivative, which is computationally costly, and reduces the training time compared to, for instance, the DGM of Sirignano and Spiliopoulos [37]; see e.g. Georgoulis et al. [10, Sec. 5].
@@
where $\mathcal{L}$ is a self-adjoint, linear operator and $F$ is a (linear) operator from $\mathcal{V}$ to $\mathcal{V}^{*}$. This PDE is then discretized using, for example, the backward Euler differentiation scheme, which yields
@@
Let us mention that the convergence results hold for certain non-linear operators $F$, assuming that the PDE is well-posed. Next, we present examples of PDEs that have been treated by DGFMs, and their applications.
@@ 3.1. Well-posedness
According to Hilber et al. [15, Theorem 3.2.2], we only need to verify that the bilinear form $\left\langle\mathcal{A}u,v\right\rangle_{\mathcal{H}^{-1},\mathcal{H}_{0}^{1}}$ is continuous and satisfies the “Gårding inequality”, where
@@ 3.2. Time stepping
The proof follows directly from Theorem 2.1 in Akrivis et al. [2]. Indeed, using that $U^{0}=u\left(0\right)$, we can show by direct, but tedious, calculations that the assumptions of [2, p. 523] are satisfied for $\lambda<1$ and $q=1$. ∎
@@ 3.4. Neural network approximation and a version of the Universal Approximation Theorem
Observe that $\mathcal{C}\left(\psi\right)\subset\mathcal{H}_{0}^{1}(\mathbb{R}^{d})$, since $\psi\in C^{\infty}_{c}(\mathbb{R}^{d})$. Assume that $\mathcal{C}\left(\psi\right)$ is not dense in $\mathcal{H}_{0}^{1}(\mathbb{R}^{d})$ then, as a corollary of the Hahn–Banach extension theorem, see e.g. van Neerven [39, Corollary 4.12], there exists a non-zero continuous linear functional $G$ on $\mathcal{H}_{0}^{1}(\mathbb{R}^{d})$ such that for any $\zeta\in\mathcal{C}\left(\psi\right)$,
@@
Since $g^{\varepsilon}_{1}-\nabla\cdot\left(g^{\varepsilon}_{2}\right)$ is continuous, see e.g. van Neerven [39, Proposition 11.1], by Lemma 3.12, we conclude that $g^{\varepsilon}_{1}-\nabla\cdot\left(g^{\varepsilon}_{2}\right)=0$. Then, for any $f\in C^{\infty}_{c}(\mathbb{R}^{d})$,
@@ 3.5. Convergence of the minimizer
Next, we prove that $I^{k}\left(w_{m}\right)\to I^{k}\left(w_{*}\right)$ implies that $w_{m}\to w_{*}$ in $\mathcal{H}_{0}^{1}$. Let us first notice that $w_{m}\rightharpoonup w_{*}$. Otherwise, there exists a subsequence $(w_{m_{i}})$, an ${\varepsilon}>0$ and a nonzero functional $f$ such that $\left|f\left[w_{m_{i}}\right]-f\left[w_{*}\right]\right|\geq{\varepsilon}$. Since $(w_{m_{i}})$ is bounded in $\mathcal{H}_{0}^{1}$ (otherwise $I^{k}\left(w_{m_{i}}\right)$ is unbounded), it is pre-weakly compact (which means it has a weakly convergent subsequence, see e.g. van Neerven [39, Corollary 4.56]). Let us denote one of its weak limits by $w_{**}$. Using Lemma 3.5, $I^{k}\left(w_{*}\right)=\lim_{i\to\infty}I^{k}\left(w_{m_{i}}\right)\geq I^{k}\left(w_{**}\right)$. This inequality implies $w_{*}=w_{**}$ by the uniqueness of the minimizer. Hence $w_{m}\rightharpoonup w_{*}$, a contradiction with $\left|f\left[w_{m_{i}}\right]-f\left[w_{*}\right]\right|\geq{\varepsilon}$. Therefore, $w_{m}\rightharpoonup w_{*}$.
@@ 4.1. Convergence of the trained neural network
for some $r_{n}$ increasing with $n$. We restrict the domain of the parameters $(\beta^{i},\alpha^{i},c^{i})$ to $r_{n},r_{n}]$ which converges to $\mathbb{R}$ as $n\to\infty$, and for $\alpha^{i}$ we also need to subtract the ball $\left(-\frac{1}{r_{n}},\frac{1}{r_{n}}\right)$. Gradient clipping is in accordance with deep learning literature, see, for example, Zhang et al. [41] and Goodfellow et al. [12, Ch. 10 and 11].
@@ 4.2. Long time behavior of the gradient flow
Using van Neerven [39, Proposition 14.13], every trace class operator is compact and positive definite. Therefore, we can do a spectral decomposition for the operator $\widetilde{\mathcal{T}}$. There exists an orthogonal basis $\{\tilde{\mathrm{e}}_{i}\}^{\infty}_{i=1}$, such that